OpenStudy (anonymous):

If a cube has a surface area of 6x^2+60+150 what is the length of a side?

jhonyy9 (jhonyy9):

6(x^2 +10x+25) - i think you have missed one x there next 60 ,yes ?

OpenStudy (anonymous):

yes I did, thanks after 60

jhonyy9 (jhonyy9):

ok - than 6(x^2 +10x +25) =6(x+5)^2 -so can you continue it ?

jhonyy9 (jhonyy9):

- so a cube how many side has ?

OpenStudy (anonymous):

x+5

jhonyy9 (jhonyy9):

answer me the question please

OpenStudy (anonymous):

The question is Find the length of a side?

jhonyy9 (jhonyy9):

yes but i have wrote to you from how many sides you will get the surface area of a cube ?

jhonyy9 (jhonyy9):

can tel me ?

OpenStudy (anonymous):

okay, thanks!

jhonyy9 (jhonyy9):

what ? not is ended again

jhonyy9 (jhonyy9):

so lateral there are 4 sides and one down and one up so this is total 6 yes ?

jhonyy9 (jhonyy9):

than from your equation we have got 6(x+5)^2

jhonyy9 (jhonyy9):

- so if 6*(x+5^2 is the surface area of cube so than from this result that one side has area of (x+5)^2 - yes ?

jhonyy9 (jhonyy9):

- so but you know that area of a square - because the side of a cube is a form of square - so area of a square is equal side time side what is side squared - yes ?

OpenStudy (anonymous):

Thank you so much for helping me!

jhonyy9 (jhonyy9):

- so than (x+5)^2 will be area of a side than from this resulted that one side will be x+5 - so than x=5 - because one lenght not can be never negativ

jhonyy9 (jhonyy9):

hope so much that was understandably sure right

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

yes, it was was! thanks again!